A ‘plane which is flying at 216 knots on a course of 310 degrees, passes over Dubbo at
3.40pm. How far west of Dubbo will it be at 4.15pm?
Give your answer to the nearest nautical
mile .(HINT: 1knot = 1 nautical mile /hour)

Answer by Alan3354(69443)

A ‘plane which is flying at 216 knots on a course of 310 degrees, passes over Dubbo at
3.40pm. How far west of Dubbo will it be at 4.15pm?
--------------
t = 35 minutes = 7/12 hours
---
d = 216*(7/12) = 126 nm from Dubbo
Distance West = 126*cos(40)